# Diff-Instruct*: Towards Human-Preferred One-step Text-to-image Generative Models

### v2

@@ 3.3 Two Ways of Incorporating Classifier-free Guidance
       freeze $\theta$, update ${\psi}$ for $K_{TA}$ rounds using SGD by minimizing
freeze $\psi$, update $\theta$ using SGD by minimizing loss
$+\alpha_{cfg}\cdot w(t)\big{\{}\bm{s}_{ref}(\operatorname{sg}[\bm{x}_{t}]|t,\bm{c})-\bm{s}_{ref}(\operatorname{sg}[\bm{x}_{t}]|t,\bm{\varnothing})\big{\}}^{T}\bm{x}_{t}\bigg{]}\bigg{\}}$ (3.5)
@@
Algorithm 1 Diff-Instruct* for training human-preferred one-step text-to-image 1-step model.
@@
The classifier-free guidance (Ho & Salimans, 2022) (CFG) uses a modified score function of a form
$\tilde{\bm{s}}_{ref}(\bm{x}_{t},t|\bm{c})\coloneqq\bm{s}_{ref}(\bm{x}_{t},t|\bm{\varnothing})+\omega\big{\{}\bm{s}_{ref}(\bm{x}_{t},t|\bm{c})-\bm{s}_{ref}(\bm{x}_{t},t|\bm{\varnothing})\big{\}}$
to replace the original conditions score function $\bm{s}_{ref}(\bm{x}_{t},t|\bm{c})$. Using CFG for diffusion models empirically leads to better sampling quality.
As is first pointed out by Luo (2024), the classifier-free guidance is related to an implicit reward function. In this part, we derive a tractable loss function that minimizes the so-called classifier-free reward, which we use together with the explicit reward $r(\cdot,\cdot)$ in DI*.
@@
$\mathcal{L}_{cfg}(\theta)=\int_{t=0}^{T}\mathbb{E}_{\bm{z}\sim p_{z},\bm{x}_{0}=g_{\theta}(\bm{z},\bm{c})\atop\bm{x}_{t}|\bm{x}_{0}\sim p(\bm{x}_{t}|\bm{x}_{0})}w(t)\bigg{\{}\bm{s}_{ref}(\operatorname{sg}[\bm{x}_{t}]|t,\bm{c})-\bm{s}_{ref}(\operatorname{sg}[\bm{x}_{t}]|t,\bm{\varnothing})\bigg{\}}^{T}\bm{x}_{t}\mathrm{d}t$ (3.6)
@@
Figure 3: A visual comparison on Parti prompts of our introduced 2.6B DI*-SDXL-1step text-to-image generative model and 12B FLUX-dev-50step and SD3.5-large-28step diffusion models. DI*-SDXL-1step model outperforms FLUX-dev-50step significantly in PickScore, ImageReward, and CLIPScore, while only taking 1.88% inference time and 29.30% GPU memory cost. Benefiting from RLHF using human feedback reward, DI*-SDXL-1step shows better aesthetic details, improved layout, and better reality than FLUX-dev-50step and SD3.5-large-28step. Prompts from top to bottom are “the skyline of New York City”, “a shiba inu”, “Pirate ship sailing into a bioluminescence sea with a galaxy in the sky, epic, 4k, ultra”, and “a volcano exploding next to a marina”.
Choice 1: Incorporating CFG via Implicit CFG Reward. Based on Theorem 3.2, we can incorporate classifier-free guidance by adding the pseudo reward (3.7) to explicit human reward to get a mixed reward. The guidance scale $\alpha_{cfg}$ and $\alpha_{rew}$ balances the strength of explicit human reward and implicit CFG reward. Choice 2: Incorporate CFG via CFG-enhanced Reference Diffusion. Besides the implicit CFG reward, we can also inject the CFG mechanism into Diff-Instruct* by replacing the naive reference diffusion with a CFG-enhanced reference diffusion which writes $\tilde{\bm{s}}_{ref}(\bm{x}_{t}|t,\bm{c})\coloneqq\bm{s}_{ref}(\bm{x}_{t}|t,\bm{c})+\alpha_{cfg}(\bm{s}_{ref}(\bm{x}_{t}|t,\bm{\varnothing}))$. This approach is straightforward. In practice, we find empirically that the choice 2 is easier to tune and results in better performances. However, choice 1 also leads to solid 1-step models. Please check out ablation studies in Table 2 for details.
@@ 3.4 The Practical Algorithm
Now it is time for us to introduce the practical algorithm. As Algorithm 1 shows, the DI* involves three models, with one 1-step model $g_{\theta}$, one reference diffusion model $\bm{s}_{ref}$ and one assistant diffusion model $\bm{s}_{\psi}$. The reference diffusion does not need to be trained, while the 1-step model and the assistant diffusion are updated alternatively. Two hyper-parameters, the $\alpha_{rew}$ and $\alpha_{cfg}$ control the strength of the explicit reward and the implicit CFG reward during training. The explicit reward model can either be an off-the-shelf reward model, such as PickScore (Kirstain et al., 2023), CLIP similarity score (Radford et al., 2021), or reward trained in-house with internal feedback data.
@@
Clearly, various choices of distance function $\mathbf{d}(.)$ result in different training algorithms. For instance, $\mathbf{d}(\bm{y}_{t})=\|\bm{y}_{t}\|_{2}^{2}$ is a naive choice. Interestingly, such a distance function has been studied in pure diffusion distillation literature in Luo et al. (2024c); Zhou et al. (2024b; a). In this paper, we draw inspiration from (Luo et al., 2024c) and find that using the so-called pseudo-Huber distance leads to better performance. The distance and corresponding loss writes $\bm{d}(\bm{y})\coloneqq\sqrt{\|\bm{y}_{t}\|_{2}^{2}+c^{2}}-c$, and $\bm{y}_{t}\coloneqq\bm{s}_{p_{\operatorname{sg}[\theta],t}}(\bm{x}_{t})-\bm{s}_{q_{t}}(\bm{x}_{t})$, then
$\mathbf{D}^{[0,T]}(p_{\theta},p_{ref})=-\bigg{\{}\frac{\bm{y}_{t}}{\sqrt{\|\bm{y}_{t}\|_{2}^{2}+c^{2}}}\bigg{\}}^{T}\bigg{\{}\bm{s}_{\psi}(\bm{x}_{t},t)-\nabla_{\bm{x}_{t}}\log p_{t}(\bm{x}_{t}|\bm{x}_{0})\bigg{\}}.$ (3.8)
Figure 4: Ablation comparisons of DI*-1step models and other SDXL-based models in Table 2.
@@ 4 Related Works
Diff-Instruct* is inspired by research on diffusion distillation (Luo, 2023) which aims to minimize distribution divergence to train one-step text-to-image models. Luo et al. (2024b) first study the diffusion distillation by minimizing the Integral KL divergence. Yin et al. (2023) generalize Diff-Instruct and add a data regression loss for better performance. Zhou et al. (2024b) study the distillation by minimizing the Fisher divergence. Luo et al. (2024c) study the distillation using the general score-based divergence. Many other works also introduced additional techniques and improved the performance (Geng et al., 2023; Kim et al., 2023; Song et al., 2023; Song & Dhariwal, ; Nguyen & Tran, 2023; Song et al., 2024; Yin et al., 2024; Zhou et al., 2024a; Heek et al., 2024; Xie et al., 2024; Salimans et al., 2024; Geng et al., 2024; Meng et al., 2022; Sauer et al., 2023b; Luo et al., 2023a; Liu et al., 2023; Gu et al., 2023; Xu et al., 2024; Ren et al., 2024; Lin et al., 2024; Zheng et al., 2024; Li et al., 2024; Berthelot et al., 2023; Zheng et al., 2022).
@@
In recent years, many works have emerged trying to align diffusion models with human preferences. There are three main lines of alignment methods for diffusion models. 1) The first kind of method fine-tunes the diffusion model over a specifically curated image-prompt dataset (Dai et al., 2023; Podell et al., 2023). 2) the second line of methods tries to maximize some reward functions either through the multi-step diffusion generation output (Prabhudesai et al., 2023; Clark et al., 2023; Lee et al., 2023) or through policy gradient-based RL approaches (Fan et al., 2024; Black et al., 2023). For these methods, the backpropagation through the multi-step diffusion generation output is expensive and hard to scale. 3) the third line, such as Diffusion-DPO (Wallace et al., 2024), Diffusion-KTO (Yang et al., 2024), tries to directly improve the diffusion model’s human preference property with raw collected data instead of reward functions. Besides the human preference alignment of diffusion models, Diff-Instruct++(Luo, 2024) recently arose as the first attempt to improve human preferences for one-step text-to-image models. Though we get inspiration from Diff-Instruct++, our Diff-Instruct* uses score-based divergences which are technically different from the KL divergence used in DI++. Besides, as we show in Section 5.1, Diff-Instruct* archives better performances than Diff-Instruct++.
Table 1: Quantitative comparisons of $1024\times 1024$ resolution leading text-to-image models on Parti(Yu et al., 2022) Prompts (the upper part) and MSCOCO-2017 validation prompts (the under part). DI*: short for Diff-Instruct*. SIM: short for Score Implicit Matching. : our implementation.
@@
DI*-SDXL (Ours, Longer Training) 1 UNet 2.6B 1.140 5.83 0.2331 32.75 2.22 Sec
@@
DI*-In-SDXL ((Human Reward + CFG)) 1 2.6B 1.031 5.69 0.2274 32.87 $(1,7.5)$
@@ 5 Experiments
We use the open-sourced SDXL of 1024$\times$1024 resolution as our reference diffusion in Algorithm 1. We construct the 1-step model with the same architecture as the SDXL model and initialize the 1-step model weights with DMD2-SDXL-1step model (Yin et al., 2024), which is a good 1-step diffusion distillation model using a combination of KL divergence and GAN loss. We use the prompts of the LAION-AESTHETIC dataset with an aesthetic score larger than 6.25, which resulting a total of 3M text prompts. We use the PickScore (Kirstain et al., 2023)(a high-quality off-the-shelf reward model trained using human feedback data) as the explicit human reward. In Table 2, we do an ablation study that compares 1-step DI* models with different guidance strategies and reward scales and finds that using inner guidance with a scale of 7.5 and an explicit human reward with a scale of 100 results in the best performances.
@@
We compare the DI*-SDXL-1step model with other leading open-sourced models of a resolution of $1024\times 1024$ on the Parti prompt benchmark, COCO-2014-30K benchmark, and also Human preference scores. These models include diffusion-based (or flow-matching) models such as FLUX-dev (flu, ), Stable-Diffusion-3.5-Large(sd, 3), and SDXL (Podell et al., 2023) with its DPO variant (Wallace et al., 2024); 1024 resolution 1-step model such as DMD2(Yin et al., 2024), Diff-Instruct(Luo et al., 2024b), Score Implicit Matching (SIM) (Luo et al., 2024c), and also Diff-Instruct++ (Luo, 2024). On Parti and COCO prompt datasets, we compute four standard scores in Table 1: the Image Reward (Xu et al., 2023), the Aesthetic Score (Schuhmann, 2022), the PickScore(Kirstain et al., 2023), and the CLIP score(Radford et al., 2021). We also evaluate the 30K-FID (Heusel et al., 2017) on the COCO validation dataset. On Human Preference Score (Wu et al., 2023) benchmarks, we calculate HPSv2.1 scores.
@@
1Step-DI*-SDXL(ours, longer training) 33.22 31.67 31.25 28.6228.62{28.62}28.62 31.19
@@ 5.1 Performances and Findings
The FLUX-dev-50step model(flu, ) is regarded as a leading open-sourced diffusion model that is trained over millions of high-quality internal data. As Table 1 shows, our best 2.6B DI*-SDXL-1step model slightly outperforms the 12B FLUX-dev-50step diffusion model and 8B SD3.5-large-28step with only 1.88% inference time and 29.3% GPU memory costs on the Parti prompt dataset which consists of challenging prompts with rich concepts. The 2.6B DI*-SDXL-1step is on par with FLUX and SD3.5 on COCO prompts in terms of preference scores and FIDs. As Table 3 shows, the DI*-SDXL-1step model has a record-breaking HPSv2.1 score of 31.19, outperforming the rest of text-to-image models of $1024\times 1024$ resolution.
@@
In Table 2, we conduct a comprehensive ablation study on the Parti prompt dataset to compare DI* (RLHF using score-based divergence) with different guidance strategies and different reward scales (which sometimes recover previous distillation approaches such as Diff-Instruct(Luo et al., 2024b) and Score Implicit Matching(Luo et al., 2024c)), and DI++(Luo, 2024)(RLHF using KL divergence). The results show that DI* using score-based divergence outperforms the DI++ model under the same configurations and reward scales. Besides, with proper reward scales, DI* models show dominant human preference scores, indicating its simple yet effective contribution to aligning one-step text-to-image models with human feedback. Besides, we also find that DI++ using KL divergence is very easy to collapse to painting-like images with over-saturated color and lighting (please refer to Figure 1 for visualizations) that have high human reward but lack diversities. Instead, using score-based divergences empirically maintains diversity while still steadily improving human preferences. This means DI* can improve images’ human preference regardless they are painting-like, photo-like, or animations. However, we do acknowledge an interesting finding of the trade-off between preference scores and CLIP scores: we find that the larger-scale explicit human reward is assigned, the worse the CLIPS score will be. We think this phenomenon reveals a natural contradiction between human-preferred data distributions and objective data distribution. The RLHF encourages the 1-step model to generate human-preferred images, which inevitably changes the underlying distribution from the ground-truth data distribution. We also find that though FLUX shows very strong human preference performances, its objective scores such as the CLIPScore and the FID are not that good even compared with smaller models.
@@
Figure 2 shows some 1024$\times$1024 images generated by DI*-SDXL-1step models. Figure 3 shows a qualitative comparison of the DI*-SDXL-1step model against FLUX and SD3.5 diffusion models. We are surprised to find that the DI*-SDXL-1step model shows better aesthetic details, improved layouts, and reality lighting than the other two models due to the RLHF. Figure 4 shows the visual comparison of DI* (with different ablative reward scales) against our initial model (the DMD2-1step model(Yin et al., 2024)), other 1-step models, and SDXL with and without DPO. While SIM (Luo et al., 2024c), DI++(Luo, 2024), and Diff-Instruct(Luo et al., 2024b) can output high-quality samples, they seem to suffer from over-saturated issues, meaning that the images have over-saturated colors and lighting that might make users feel uncomfortable. On the contrary, the DI*-1step model shows very gentle lighting and aesthetic colors, which align better with human preferences. We empirically conclude that DI* prevents the model from collapsing into high-reward painting-like images in which DI++ collapsed. Besides, the SDXL and SDXL-DPO are easy to generate painting-like images instead of photo-realistic images.

### v3

@@ 3.3 Understanding Classifier-free Guidance
       freeze $\theta$, update ${\psi}$ for $K_{TA}$ rounds by
  1. 1.
    sample prompt $\bm{c}\sim\mathcal{C}$; sample time $t\sim\pi(t)$; sample $\bm{z}\sim p_{z}(\bm{z})$;
  2. 2.
    generate fake data: $\bm{x}_{0}=\operatorname{sg}[g_{\theta}(\bm{z},\bm{c})]$; sample noisy data: $\bm{x}_{t}\sim p_{t}(\bm{x}_{t}|\bm{x}_{0})$;
  3. 3.
    update $\psi$ by minimizing loss: $\mathcal{L}(\psi)=\lambda(t)\|\bm{s}_{\psi}(\bm{x}_{t}|t,\bm{c})-\nabla_{\bm{x}_{t}}\log p_{t}(\bm{x}_{t}|\bm{x}_{0})\|_{2}^{2}$;
freeze $\psi$, update $\theta$ using SGD:
  1. 1.
    sample prompt $\bm{c}\sim\mathcal{C}$; sample time $t\sim\pi(t)$; sample $\bm{z}\sim p_{z}(\bm{z})$;
  2. 2.
    generate fake data: $\bm{x}_{0}=g_{\theta}(\bm{z},\bm{c})$; sample noisy data: $\bm{x}_{t}\sim p_{t}(\bm{x}_{t}|\bm{x}_{0})$;
  3. 3.
    explicit reward: $\mathcal{L}_{rew}(\theta)=-\alpha_{rew}r(\bm{x}_{0},\bm{c})$;
  4. 4.
    CFG reward: $\mathcal{L}_{cfg}(\theta)=\alpha_{cfg}\cdot w(t)\big{\{}\bm{s}_{ref}(\operatorname{sg}[\bm{x}_{t}]|t,\bm{c})-\bm{s}_{ref}(\operatorname{sg}[\bm{x}_{t}]|t,\bm{\varnothing})\big{\}}^{T}\bm{x}_{t}$;
  5. 5.
    score-regularization: $\mathcal{L}_{reg}(\theta)=-w(t)\big{\{}\mathbf{d}^{\prime}(\bm{s}_{\psi}(\bm{x}_{t}|t,\bm{c})-\bm{s}_{ref}(\bm{x}_{t}|t,\bm{c}))\big{\}}^{T}\big{\{}\bm{s}_{\psi}(\bm{x}_{t}|t,\bm{c})-\nabla_{\bm{x}_{t}}\log p_{t}(\bm{x}_{t}|\bm{x}_{0})\big{\}}$;
  6. 6.
    update $\theta$ by minimizing DI* loss: $\mathcal{L}_{DI*}(\theta)=\mathcal{L}_{rew}(\theta)+\mathcal{L}_{cfg}(\theta)+\mathcal{L}_{reg}(\theta)$;
@@
Algorithm 1 Diff-Instruct* Pseudo Code.
@@
The classifier-free guidance (Ho & Salimans, 2022) (CFG) uses a modified score function of the form
$\tilde{\bm{s}}_{ref}(\bm{x}_{t},t|\bm{c})\coloneqq\bm{s}_{ref}(\bm{x}_{t},t|\bm{\varnothing})+\omega\big{\{}\bm{s}_{ref}(\bm{x}_{t},t|\bm{c})-\bm{s}_{ref}(\bm{x}_{t},t|\bm{\varnothing})\big{\}},$
to replace the original conditional score function $\bm{s}_{ref}(\bm{x}_{t},t|\bm{c})$. Using CFG for diffusion models empirically leads to better sampling fidelity but with a cost of diversity.
As is first pointed out by Luo (2024), the classifier-free guidance is related to an implicit RLHF. In this part, we derive a tractable loss function that minimizes the so-called classifier-free reward, which we use together with the explicit reward $r(\cdot,\cdot)$ in DI*.
@@
$\mathcal{L}_{cfg}(\theta)=\int_{0}^{T}\mathbb{E}_{\bm{z}\sim p_{z},\bm{x}_{0}=g_{\theta}(\bm{z},\bm{c}),\atop\bm{x}_{t}\sim p(\bm{x}_{t}|\bm{x}_{0})}w(t)\big{[}\bm{s}_{ref}(\operatorname{sg}[\bm{x}_{t}]|t,\bm{c})$ (3.5)
$-\bm{s}_{ref}(\operatorname{sg}[\bm{x}_{t}]|t,\bm{\varnothing})\big{]}^{T}\bm{x}_{t}\mathrm{d}t.$
@@
Choice 1: Incorporating CFG via Implicit CFG Reward. Based on Theorem 3.2, we can incorporate classifier-free guidance by adding the pseudo reward (3.6) to explicit human reward to get a mixed reward. The guidance scale $\alpha_{cfg}$ and $\alpha_{rew}$ balances the strength of explicit human reward and implicit CFG reward.
Choice 2: Incorporate CFG via CFG-enhanced Reference Diffusion. Besides the implicit CFG reward, we can also inject the CFG mechanism into Diff-Instruct* by replacing the naive reference diffusion with a CFG-enhanced reference diffusion which writes $\tilde{\bm{s}}_{ref}(\bm{x}_{t}|t,\bm{c})\coloneqq\bm{s}_{ref}(\bm{x}_{t}|t,\bm{c})+\alpha_{cfg}(\bm{s}_{ref}(\bm{x}_{t}|t,\bm{\varnothing}))$. This approach is straightforward. In practice, we find empirically that the choice 2 is easier to tune and results in better performances. However, choice 1 also leads to solid one-step models. Please check out ablation studies in Table 2 for details.
@@ 3.4 The Practical Algorithm
As Algorithm 1 shows, the DI* involves three models, with one one-step model $g_{\theta}$, one reference diffusion model $\bm{s}_{ref}$ and one assistant diffusion model $\bm{s}_{\psi}$. The reference diffusion is frozen, while the one-step model and the assistant diffusion are updated alternately. Two hyperparameters, the $\alpha_{rew}$ and $\alpha_{cfg}$, control the strength of the explicit reward and the implicit CFG reward during training. The explicit rewards can either be off-the-shelf reward models, such as PickScore (Kirstain et al., 2023), CLIP score (Radford et al., 2021), or rewards trained in-house with internal feedback data.
@@
Various choices of distance function $\mathbf{d}(.)$ result in different training algorithms. For instance, $\mathbf{d}(\bm{y}_{t})=\|\bm{y}_{t}\|_{2}^{2}$ is a naive choice. This distance function has been studied in the pure diffusion distillation literature (Luo et al., 2024c; Zhou et al., 2024b, a). In this paper, we draw inspiration from (Luo et al., 2024c) and find that using the so-called pseudo-Huber distance leads to better performance. The distance writes $\bm{d}(\bm{y})\coloneqq\sqrt{\|\bm{y}_{t}\|_{2}^{2}+c^{2}}-c$, and $\bm{y}_{t}\coloneqq\bm{s}_{p_{\operatorname{sg}[\theta],t}}(\bm{x}_{t})-\bm{s}_{q_{t}}(\bm{x}_{t})$, then
$\mathbf{D}^{[0,T]}(p_{\theta},p_{ref})$ (3.7)
=$=-\bigg{\{}\frac{\bm{y}_{t}}{\sqrt{\|\bm{y}_{t}\|_{2}^{2}+c^{2}}}\bigg{\}}^{T}\bigg{\{}\bm{s}_{\psi}(\bm{x}_{t},t)-\nabla_{\bm{x}_{t}}\log p_{t}(\bm{x}_{t}|\bm{x}_{0})\bigg{\}}.$
@@ 4 Related Works
Diff-Instruct* is inspired by research on diffusion distillation (Luo, 2023), which aims to minimize distribution divergence to train one-step text-to-image models. Luo et al. (2024b) first studies the diffusion distillation by minimizing the Integral KL divergence. Yin et al. (2023) generalizes Diff-Instruct and adds a data regression loss for better performance. Zhou et al. (2024b) study the distillation by minimizing the Fisher divergence. Luo et al. (2024c) study the distillation using the general score-based divergence. Many other works also introduced additional techniques and improved the performance (Geng et al., 2023; Kim et al., 2023; Song et al., 2023; Song & Dhariwal, 2024; Nguyen & Tran, 2023; Song et al., 2024; Yin et al., 2024; Zhou et al., 2024a; Heek et al., 2024; Xie et al., 2024; Salimans et al., 2024; Geng et al., 2024; Meng et al., 2022; Sauer et al., 2023b; Luo et al., 2023a; Liu et al., 2023; Gu et al., 2023; Xu et al., 2024; Ren et al., 2024; Lin et al., 2024; Zheng et al., 2024; Li et al., 2024; Berthelot et al., 2023; Zheng et al., 2022).
@@
In recent years, many works have emerged trying to align diffusion models with human preferences. There are three main lines of alignment methods for diffusion models. 1) The first kind of method fine-tunes the diffusion model over a specifically curated image-prompt dataset (Dai et al., 2023; Podell et al., 2023). 2) The second line of methods tries to maximize some reward functions either through the multi-step diffusion generation output (Prabhudesai et al., 2023; Clark et al., 2023; Lee et al., 2023) or through policy gradient-based RL approaches (Fan et al., 2024; Black et al., 2023). For these methods, the backpropagation through the multi-step diffusion generation output is expensive and hard to scale. 3) The third line, such as Diffusion-DPO (Wallace et al., 2024), Diffusion-KTO (Yang et al., 2024), and MaPO (Hong et al., 2024) tries to directly improve the diffusion model’s human preference property with raw collected data instead of reward functions. Besides the human preference alignment of diffusion models, Diff-Instruct++(Luo, 2024) recently arose as the first attempt to improve human preferences for one-step text-to-image models. Though we get inspiration from Diff-Instruct++, our Diff-Instruct* uses score-based divergences, which are technically different from the KL divergence used in DI++. Besides, as we show in Section 5.1, Diff-Instruct* outperforms Diff-Instruct++.
Table 1: Quantitative comparisons of $1024\times 1024$ resolution leading text-to-image models on Parti(Yu et al., 2022) Prompts (the upper part) and MSCOCO-2014 validation prompts (the under part). DI*: short for Diff-Instruct*. SIM: short for Score Implicit Matching. : our implementation.
@@
DI*-SDXL (Ours, Long Train) 1 UNet 2.6B 1.140 5.83 0.2331 32.75 2.22 Sec
DI*-SDXL-DDPO (Ours, Long Train) 1 UNet 2.6B 1.210 5.90 0.2342 32.91 2.22 Sec
@@
DI*-In-SDXL (Human Reward + CFG)) 1 2.6B 1.031 5.69 0.2274 32.87 $(1,7.5)$
@@ 5 Experiments
We use the open-sourced SDXL of 1024$\times$1024 resolution as our reference diffusion in Algorithm 1. We construct the one-step model with the same architecture as the SDXL model and initialize the one-step model weights with the DMD2-SDXL-1step model (Yin et al., 2024), which is a good one-step text-to-image model distilled by using a combination of Diff-Instruct loss and GAN loss. We use the prompts of the LAION-AESTHETIC dataset with an aesthetic score larger than 6.25, resulting in a total of 2.2 million text prompts. We use the PickScore (Kirstain et al., 2023)(a high-quality off-the-shelf reward model trained using human feedback data) as the explicit human reward. In Table 2, we do an ablation study that compares one-step DI* models with different guidance strategies and reward scales and finds that using inner guidance with a scale of 7.5 and an explicit human reward with a scale of 100 results in the best performance.
@@
We compare the DI*-SDXL-1step model with other leading open-sourced models of a resolution of $1024\times 1024$ on the Parti prompt benchmark, COCO-2014-30K benchmark, and also Human preference scores. These models include diffusion-based (or flow-matching) models such as FLUX-dev (Lab, 2024), Stable-Diffusion-3.5-Large(AI, 2024), and SDXL (Podell et al., 2023) with its DPO variant (Wallace et al., 2024); 1024 resolution one-step models such as DMD2(Yin et al., 2024), Diff-Instruct(Luo et al., 2024b), Score Implicit Matching (SIM) (Luo et al., 2024c), and also Diff-Instruct++ (Luo, 2024). On Parti and COCO prompt datasets, we compute four standard scores in Table 1: the Image Reward (Xu et al., 2023), the Aesthetic Score (Schuhmann, 2022), the PickScore (Kirstain et al., 2023), and the CLIP score (Radford et al., 2021). We also evaluate the 30K-FID (Heusel et al., 2017) on the COCO validation dataset. On Human Preference Score (Wu et al., 2023) benchmarks, we calculate HPSv2.1 scores.
@@
1Step-DI*-SDXL(ours, longer training) 33.2233.22{33.22}33.22 31.6731.67{31.67}31.67 31.2531.25{31.25}31.25 28.6228.62{28.62}28.62 31.1931.19{31.19}31.19
1Step-DI*-SDXL-DDPO(ours, longer training) 33.92 32.80 32.71 29.62 32.26
@@ 5.1 Performances and Findings
The FLUX-dev-50step model (Lab, 2024) is recognized as a leading open-sourced diffusion model that is trained on millions of high-quality internal data samples. As shown in Table 1, our best 2.6B DI*-SDXL-1step model outperforms the 12B FLUX-dev-50step diffusion model and 8B SD3.5-large-28step using only 1.88% inference time and 29.3% GPU memory costs on the Parti prompt dataset which consists of challenging prompts with rich concepts. The 2.6B DI*-SDXL-1step is on par with FLUX and SD3.5 on COCO prompts in terms of preference scores and FIDs. As Table 3 shows, the DI*-SDXL-1step model has a record-breaking HPSv2.1 score of 31.19, outperforming the rest of the text-to-image models of $1024\times 1024$ resolution. These empirical results confirm that proper post-training enables small one-step models to outperform huge multi-step models, analogous to the story of David and Goliath. See Figure 2 for visual comparisons with the 12B FLUX-Dev and the 8B Stable Diffusion 3.5-Large models.
@@
In Table 2, we conduct a comprehensive ablation study on the Parti prompt dataset to compare DI* (RLHF using score-based divergence) with different guidance strategies and reward scales (which recover previous distillation approaches such as Diff-Instruct(Luo et al., 2024b) and Score Implicit Matching(Luo et al., 2024c)), and DI++(Luo, 2024)(RLHF with KL divergence).
The results show that DI* using score-based divergence outperforms the DI++ model under the same configurations and reward scales. Notably, with proper reward scales, DI* models demonstrate superior human preference scores, indicating its simple yet effective contribution to aligning one-step text-to-image models with human feedback. Critically, DI++ with KL divergence tends to collapse to painting-like images with over-saturated color and lighting (Please refer to Figure 1 for visualizations), which lack diversity despite high rewards. Instead, score-based divergences empirically maintain diversity while steadily improving human preferences, suggesting that DI* can improve images’ human preference scores regardless of their styles in paintings, photos, or animations.
In addition, we acknowledge the trade-off between preference scores and CLIP scores: increasing explicit human reward scales degrades CLIP scores. This phenomenon reveals a natural contradiction between human-preferred and objective data distributions. RLHF encourages the one-step model to generate human-preferred images, which inevitably shifts the underlying distribution from the ground-truth data distribution. This aligns with observations that FLUX, despite strong preference scores, underperforms in CLIPScore and FID compared to smaller models.
DI* is compatible with Pre-aligned Diffusion Models
Beyond the ablation comparisons of different explicit reward strengths and CFG strengths, we also find that DI* is compatible with pre-aligned reference diffusion models. As Table 1 and Table 3 show, if we replace the reference diffusion model from SDXL to open-sourced SDXL-DDPO (Stable Diffusion XL with Diffusion Direct Preference Optimization (Wallace et al., 2024)), the preference alignment effects of the resulting one-step model are measured by PickScore, ImageReward, AestheticScore, and HPSv2.1, and were further improved. This demonstrates the compatibility and robustness of Diff-Instruct* with different reference diffusion models. In practice, we recommend that users use pre-aligned diffusion models that
@@
Figure 1 shows some 1024$\times$1024 images generated by DI*-SDXL-1step models. Figure 2 shows a qualitative comparison of the DI*-SDXL-1step model against FLUX and SD3.5 diffusion models. The DI*-SDXL-1step model shows better aesthetic details, improved layouts, and reality lighting compared to FLUX and SD3.5, attributed to our score post-training. Figure 4 shows the visual comparison of DI* (with different ablative reward scales) against our initial model (the DMD2-1step model(Yin et al., 2024)), other one-step models, and SDXL with and without DPO. While SIM (Luo et al., 2024c), DI++(Luo, 2024), and Diff-Instruct(Luo et al., 2024b) can output high-quality samples, they suffer from oversaturation issues, meaning that the images have over-saturated colors and lighting that might make users feel uncomfortable. On the contrary, the DI*-1step model shows very gentle lighting and aesthetic colors, which align better with human preferences. We empirically conclude that DI* prevents the model from collapsing into high-reward painting-like images, in which DI++ collapsed. Besides, the SDXL and SDXL-DPO are easy to generate painting-like images instead of photo-realistic images.